OpenStudy (anonymous):

Select one of the factors of x3 + 3x2 + 7x + 21.

OpenStudy (callisto):

x^3 + 3x^2 + 7x + 21 Take out the common factor = x^2 (x+3) +7(x+3) Take out the common factor again ... =(x+3) (x^2+7)

OpenStudy (anonymous):

(x + 7) (x2 + 3) (x + 21) None of the above So would i choose none of the above because none of the answer choices show your answer?

OpenStudy (callisto):

Hmm.. actually what do you think for your own answer, not my answer?

OpenStudy (callisto):

If i were you , i would choose none of the above. But you need to know that, to do such kind of questions, you'd better do the factorization, or use the remainder theorem to check it.